SLAC PEP-II
BABAR
SLAC<->RAL
Babar logo
HEPIC E,S & H Databases PDG HEP preprints
Organization Detector Computing Physics Documentation
Personnel Glossary Sitemap Search Hypernews
Unwrap page!
Det. Search
Who's who?
Meetings
FAQ
Images
Archive
Systems
Performance
Intern. region
Vertex Tracker
Drift chamber
DIRC
Calorimeter
IFR
LST
Magnet
Electronics
Trigger
Operations
Run Coordination
Contact Experts
Shift Takers Info
Operations Manual
Electronic Logbook
Ops Hypernews
Shift Signup
Check this page for HTML 4.01 Transitional compliance with the
W3C Validator
(More checks...)

Minutes of the Trigger Operation Meeting (Jan.31, 2006)

Attendees: Kris Chaisanguanthum, Al Eisner, Fang Fang, Arik Kreisel, Selina Li,Rahmat Rahmat, Dong Su, Dan Walker

Operations

  • BaBar is taking low lumi physics runs. There were quite some periods of non-data-taking time here and there for the past week.
  • The Trigger is working okay; however there were a lot of out-of-frame messages seen in the ROM logfile as well as TSF damages showed up in X15 in particular. See the DCT/DCZ section below for more information.
  • The shifters saw a lot of DCT-GLT and EMT-GLT Transmission errors in run 61706 due to the L1 trigger rate of 3.4 kHz. The rate of this type of errors is expected to grow with L1 rate. However, this high trigger rate is not expected at a luminosity of just 3.7e33. It turned out that the background was terrible and was evident in other detector subsystems as well.

EMT

  • Dan showed the new plots for fast monitoring .ps archives: Page 11 and 12 of the fastmon plots for run 61798 and run 61801. It was suggested to plot the raw data as well in the upper right hand corner of the plots in Page 11 for the EMT prescale monitoring.
  • Dan also showed the EMT deadtime and the total deadtime for Jan.27 to Jan.31. It might worth trying coarser binning (say, by day) to avoid large fluctuation from bin to bin. Su Dong pointed out that the deadtime depends on L1 rate, so it would be helpful to plot EMT deadtime vs L1 rate.
  • After we determine a good EMT prescale factor to use, we would like to have only 2 physics keys eventually: one is physics, and the other for physics with the EMT prescale.

DCT/DCZ

GLT

  • Last Wednesday when Selina gave the monthly trigger report after the detector operation meeting, there was a question about possibility of further reducing the deadtime due to GLT. Su Dong pointed out that it doesn't make sense to prescale the GLT data as it's not a FEX problem. Valerie was about to work on sparsifying the GLT a while ago, but it seemed to have no gain so it was abandoned. If we were to reduce the Glink data, the firmware issue would come up.

Online

  • The new DCT online lib was successfully tested with cosmics and DCH-DCT calibration. Fang confirmed with Matt Weaver that we ar enot using batching, so we don't need a large space for the poolsize. Fang is now working on getting the online codes checked in.

L3

  • Nothing to report for the L3 operation.
  • Arik showed some plots such as efficiency and purity for L3 tracking, comparing the new and old tracking code (blue is new code, and red is old). Basically, the change is just by doing more accurate derivative for L3 track fitting. It was suggested that Arik runs OPR with the new code to see if there is any difference in efficiency as well as other quantities that we normally monitor in OPR. Su Dong also suggested Arik to make the efficiency and purity plots with cuts applied. For example, when plotting tan(theta), he should already have cut on d0, Z0 and Pt.

Data Quality and Monitoring

  • Run 5 date checked ok.

This page is created by Selina Li (selina@slac.stanford.edu); however, you are welcome to edit and/or correct above items, specially for those who attended the meeting. Thanks.